# Discrete Approximate Circle Bundles

### v2

@@ 2.1 Theory Of Fiber Bundles
Čech Cohomology: The cocycle condition (1) and Part B of the Fiber Bundle Construction Lemma suggest a cohomological perspective to fiber bundle classification. This perspective can be articulated using the language of Čech cohomology, which we briefly describe below.
Let $\mathcal{U}$ be an open cover of a topological space $B$ and let $\mathcal{F}$ be a presheaf of abelian groups on $B$. For $k\geq 0$, define $\check{C}^{k}(\mathcal{U};\mathcal{F})$ and $\delta^{k}:\check{C}^{k}(\mathcal{U};\mathcal{F})\to\check{C}^{k+1}(\mathcal{U};\mathcal{F})$ by
$\check{C}^{k}(\mathcal{U};\mathcal{F})=\prod_{\alpha_{0}<\cdots<\alpha_{k}}\mathcal{F}(U_{\alpha_{0}}\cap\cdots\cap U_{\alpha_{k}})$
$(\delta^{k}\Omega)_{\alpha_{0}\cdots\alpha_{k+1}}=\sum_{j=0}^{k+1}(-1)^{j}\Omega_{\alpha_{0}\cdots\widehat{\alpha_{j}}\cdots\alpha_{k+1}}|_{U_{\alpha_{0}}\cap\cdots\cap U_{\alpha_{k+1}}}$
Then $(\check{C}^{*}(\mathcal{U};\mathcal{F}),\delta^{*})$ is a cochain complex, and the associated cohomology is called Čech cohomology subordinate to $\mathcal{U}$ with coefficients in $\mathcal{F}$.
Note that if $\mathcal{F}$ is a presheaf of non-abelian groups, we can still define $\check{C}^{k}(\mathcal{U};\mathcal{F})$ and $\delta^{k}$ as above, replacing negation by group inversion in the coboundary formula. In general, however, we do not have $\ker\delta^{k+1}\subseteq\text{im }\delta^{k}$ so $\check{H}^{k}(\mathcal{U};\mathcal{F})$ is not well-defined for $k>0$. For $k=1$, we still have a well-defined cohomology set $\check{H}^{1}(\mathcal{U};\mathcal{F})$, though the classes do not have a natural group structure.
Let $\mathcal{F}$ be a presheaf of groups over $B$, and let $\mathcal{U}$ be an open cover of $B$. If $\mathcal{V}$ is a refinement of $\mathcal{U}$, one has an induced map $\check{H}^{*}(\mathcal{U};\mathcal{F})\to\check{H}^{*}(\mathcal{V};\mathcal{F})$ in cohomology. The open covers of $B$ form a directed set with respect to refinements, so the corresponding Čech cohomology groups form a direct system of abelian groups. The direct limit is denoted $\check{H}^{*}(B;\mathcal{F})$ and called the Čech cohomology of $B$. If $B$ is paracompact, then for any topological group $G$, the fiber bundle construction lemma implies a one-to-one correspondence between $\check{H}^{1}(B;\mathcal{C}_{\text{Homeo}(F)})$ and $\text{Bun}_{F}(B)$, where $\mathcal{C}_{\text{Homeo}(F)}$ denotes the sheaf of continuous $\text{Homeo}(F)$-valued functions on $B$.
We state the following additional properties of Čech cohomology without proof. For further details, see [Bre97].
@@
    If $B$ has the homotopy type of a CW complex and $G$ is abelian, then $\check{H}^{*}(B;\underline{G})\cong H^{*}(B;G)$, where $\underline{G}$ denotes the sheaf of locally constant $G$-valued functions on $B$.
@@
    If $\varphi:G\to H$ is a continuous homomorphism of topological groups, there is an induced cochain map $\varphi_{*}:\check{Z}^{*}(B;\mathcal{C}_{G})\to\check{Z}^{*}(B;\mathcal{C}_{H})$ given by $(\varphi_{*}\Omega)_{\alpha_{0}\cdots\alpha_{k}}=\varphi(\Omega_{\alpha_{0}\cdots\alpha_{k}})$. Moreover, if $\varphi$ is a homotopy equivalence, then $\varphi_{*}$ induces an isomorphism in cohomology.
@@
    If $F\to G\to H$ is a central extension of topological groups, there is an induced short exact sequence $\mathcal{C}_{F}\to\mathcal{C}_{G}\to\mathcal{C}_{H}$ of sheaves, which in turn gives rise to a short exact sequence in Čech cohomology. Moreover, if $F$ is abelian, there is an induced connecting homomorphism $\Delta:\check{H}^{1}(B;\mathcal{C}_{H})\to\check{H}^{2}(B;\mathcal{C}_{F})$ which yields a long exact sequence.
@@
and $\pi^{\prime}:f^{*}E\to B^{\prime}$ is given by $\pi^{\prime}(b^{\prime},e)=b^{\prime}$. Note that if $\xi$ has fiber $F$ then so does $f^{*}\xi$. Moreover, one can show that homotopic maps yield isomorphic pullback bundles.
@@
For any paracompact space $B$, there is a bijective correspondence between $[B,BG]$ and $\text{Prin}_{G}(B)$ given by $f]\mapsto[f^{*}\xi_{G}]$, where $\xi_{G}$ is a universal $G$-bundle. Given any $\xi]\in\text{Prin}_{G}(B)$, a map $f:B\to BG$ such that $\xi\cong f^{*}\xi_{G}$ is called a classifying map for $\xi$.
Lastly, we note that the classifying space construction is functorial in the following sense: if $\varphi:G\to H$ is a continuous homomorphism of topological groups, there is an induced map $B\varphi:BG\to BH$ with the property that if $f:B\to BG$ classifies a principal $G$-bundle $\xi$, then $B\varphi\circ f:B\to BH$ classifies $\varphi^{*}\xi$.
@@ 2.2 Cohomology With Local Coefficients
We provide a brief overview of cohomology with local coefficients which we use in our classification of circle bundles. For additional details, see [Eil47], [Hat02].
@@
Given a topological space $B$ and an abelian group $\Gamma$, a local system for $B$ with coefficients in $\Gamma$ is a locally constant sheaf $\mathcal{L}$ with stalk $\Gamma$.
We will primarily be interested in the case where $\Gamma$ is a topological group and $B$ a path-connected paracompact space (in particular, $B$ has a universal cover $p:\widetilde{B}\to B$). In this setting, a local system on $B$ admits a variety of important equivalent characterizations:
@@
Suppose $G$ is a topological group and $B$ is a path-connected space with universal cover $p:\widetilde{B}\to B$. A local system on $B$ with coefficients in $G$ can be equivalently defined as any one of the following:
@@
    A sheaf $\mathcal{L}$ with stalk $G$ such that $p^{*}\mathcal{L}$ is a constant sheaf on $\widetilde{B}$,
@@
    A monodromy representation $\rho:\pi_{1}(B,b_{0})\to\text{Aut}(G)$,
@@
    A flat fiber bundle over $B$ with fiber $G$ and structure group $\text{Aut}(G)$,
@@
    A Čech cocycle $\omega\in\check{Z}^{1}(B;\underline{\text{Aut}(G)})$.
@@
If $\mathcal{L}$ is a locally-constant sheaf, then $p^{*}\mathcal{L}$ is also locally constant. Since $\widetilde{B}$ is simply-connected, $\mathcal{L}$ must be constant. Conversely, if $\mathcal{L}$ is a sheaf on $B$ and $p^{*}\mathcal{L}$ is constant, then clearly $\mathcal{L}$ is locally-constant.
1 $\to$ 2: Since the unit interval $I=[0,1]$ is contractible, the sheaf $\gamma^{*}\mathcal{L}$ on $I$ is constant for any loop $\gamma:I\to B$ based at $b_{0}$. We therefore have an induced automorphism of $G$ by identifying $(\gamma^{*}\mathcal{L})_{0}$ with $(\gamma^{*}\mathcal{L})_{1}$, and the automorphisms induced by homotopic loops coincide. We therefore have a map $\rho:\pi_{1}(B,b_{0})\to\text{Aut}(G)$, and one can easily check that this map is a representation.
2 $\to$ 3: Note that $p:\widetilde{B}\to B$ is a principal $\pi_{1}(B,b_{0})$-bundle, where $\pi_{1}(B,b_{0})$ acts on $\widetilde{B}$ by deck transformations. In particular, this bundle is flat since $\pi_{1}(B,b_{0})$ is a discrete group. A monodromy representation gives a left $\pi_{1}(B,b_{0})$-action on $G$ by automorphisms, so we can form the associated bundle $\widetilde{B}\times_{\rho}G\to B$.
@@
3 $\to$ 1: If $\xi$ is a flat fiber bundle over $B$ with fiber $G$ and structure group $\text{Aut}(G)$, the sheaf of locally-constant sections of $\xi$ is a locally-constant sheaf on $B$ with stalk $L$.
Remark 2.5.
The equivalences above all descend to the level of isomorphism classes. For example, two locally-constant sheaves on $B$ are isomorphic if and only if they induce equivalent monodromy representations.
@@
A local system with stalk $\mathbb{Z}$ is called an orientation sheaf. Given a cocycle $\omega\in Z^{1}(B;\mathbb{Z}_{2})$, we will denote the orientation sheaf associated with $\omega$ by $\mathbb{Z}_{\omega}$. In particular, if $B$ is a smooth manifold, the orientation sheaf associated with $w_{1}(TB)$ is called the orientation sheaf of the manifold.
By Proposition 2.4, we can think of an orientation sheaf over $B$ as the sheaf of locally-constant sections of the associated bundle $B_{\omega}\times_{\mathbb{Z}_{2}}\mathbb{Z}$, where $B_{\omega}$ is a double cover of $B$. More generally, given a topological group $G$ with a $\mathbb{Z}_{2}$-action, we will be interested in sheaves of sections of associated bundles of the form $B_{\omega}\times_{\mathbb{Z}_{2}}G$:
@@
$\mathcal{P}_{G,\omega}(V)=\{\text{families }\{f_{j}\in\mathcal{C}_{G}(U_{j}\cap V)\}:\ f_{j}|_{U_{j}\cap U_{k}\cap V}=\omega_{jk}\cdot f_{k}|_{U_{j}\cap U_{k}\cap V}\}$
The sheaf $\mathcal{C}_{G,\omega}$ of continuous $G$-valued functions on $B$ twisted by $\omega$ is the sheafification of $\mathcal{P}_{G,\omega}$. Note that $\mathcal{C}_{G,\omega}$ is equivalently defined as the sheaf of sections of the associated bundle $B_{\omega}\times_{\mathbb{Z}_{2}}G$, where $p_{\omega}:B_{\omega}\to B$ denotes the orientation double cover represented by $\omega$.
If $G$ is a discrete group, we will write $G_{\omega}$ instead of $\mathcal{C}_{G,\omega}$.
@@
Let $\{f_{j}\in\mathcal{C}_{G}(U_{j})\}_{j\in J}$ such that $f_{j}|_{U_{j}\cap U_{k}}=\omega_{jk}\cdot f_{k}|_{U_{j}\cap U_{k}}$ whenever $U_{j}\cap U_{k}$ is nonempty, and let $\{f^{\prime}_{j}=\phi_{j}\cdot f_{j}\in\mathcal{C}_{G}(U_{j})\}_{j\in J}$. Since $\omega=\omega^{\prime}\cdot\delta^{0}\phi$, we have
$\tilde{f}_{j}|_{U_{j}\cap U_{k}}=\phi_{j}\cdot f_{j}|_{U_{j}\cap U_{k}}=(\phi_{j}\omega_{jk})\cdot f_{k}|_{U_{j}\cap U_{k}}=\phi_{k}\cdot f_{k}|_{U_{j}\cap U_{k}}=\tilde{f}_{k}|_{U_{j}\cap U_{k}}$
@@
One can also define a notion of singular (co)homology with local coefficients (in fact, any (co)homology theory can be promoted to a theory with local coefficients). The details are not important for us, but we note that when $B$ is paracompact, the singular and Čech cohomology groups with coefficients in $\mathcal{L}$ are isomorphic. We also have the following generalization of Poincare duality:
@@
Let $B$ be a closed connected $n$-manifold, and let $\mathcal{L}$ be a local system of abelian groups. Then $H^{k}(B;\mathcal{L})\cong H_{n-k}(B;\mathcal{L}\otimes\mathcal{O})$ for all $k$, where $\mathcal{O}$ denotes the orientation sheaf of $B$.

### v3

@@ 2.1 Fiber Bundles
Čech Cohomology: The cocycle condition (1) and Part B of the Fiber Bundle Construction Lemma suggest a cohomological perspective to fiber bundle classification. This perspective can be articulated using the language of Čech cohomology, which we briefly describe below. See [18, Chapter IX] for definitions.
Let $\mathcal{U}=\{U_{\alpha}\}_{\alpha\in\Lambda}$ be an open cover of a topological space $B$ and let $\mathcal{F}$ be a presheaf of abelian groups on $B$. For $k\geq 0$, define $\check{C}^{k}(\mathcal{U};\mathcal{F})$ and $\delta^{k}:\check{C}^{k}(\mathcal{U};\mathcal{F})\to\check{C}^{k+1}(\mathcal{U};\mathcal{F})$ by
$\check{C}^{k}(\mathcal{U};\mathcal{F})=\prod_{(\alpha_{0},\ldots,\alpha_{k})}\mathcal{F}(U_{\alpha_{0}\cdots\alpha_{k}})$
$(\delta^{k}\Omega)_{\alpha_{0}\cdots\alpha_{k+1}}=\sum_{j=0}^{k+1}(-1)^{j}\Omega_{\alpha_{0}\cdots\widehat{\alpha_{j}}\cdots\alpha_{k+1}}|_{U_{\alpha_{0}\cdots\alpha_{k+1}}}$
where $U_{\alpha_{0}\cdots\alpha_{k}}=U_{\alpha_{0}}\cap\ldots\cap U_{\alpha_{k}}$ and $\alpha_{0}\cdots\widehat{\alpha_{j}}\cdots\alpha_{k+1}$ is the result of removing $\alpha_{j}$ from the original tuple. Then $(\check{C}^{*}(\mathcal{U};\mathcal{F}),\delta^{*})$ is a cochain complex, and the associated cohomology groups $\check{H}^{*}(\mathcal{U};\mathcal{F})$ are the Čech cohomology subordinate to $\mathcal{U}$ with coefficients in $\mathcal{F}$.
Note that if $\mathcal{F}$ is a presheaf of non-abelian groups, we can still define $\check{C}^{k}(\mathcal{U};\mathcal{F})$ and $\delta^{k}$ as above, replacing the minus sign by group inversion in the coboundary formula. In general, however, we do not have $\text{im }\delta^{k}\subseteq\ker\delta^{k+1}$ so $\check{H}^{k}(\mathcal{U};\mathcal{F})$ is not well-defined for $k>0$. For $k=1$, we still have a well-defined cohomology set $\check{H}^{1}(\mathcal{U};\mathcal{F})$, though it does not have a natural group structure.
Let $\mathcal{F}$ be a presheaf of groups over $B$, and let $\mathcal{U}$ be an open cover of $B$. If $\mathcal{V}$ is a refinement of $\mathcal{U}$, one has an induced map $\check{H}^{*}(\mathcal{U};\mathcal{F})\to\check{H}^{*}(\mathcal{V};\mathcal{F})$ in cohomology. The open covers of $B$ form a partially ordered set with respect to refinements, so the corresponding Čech cohomology groups (or sets) form a directed system (of groups or sets, respectively). The direct limit is denoted $\check{H}^{*}(B;\mathcal{F})$ and called the Čech cohomology of $B$ with coefficients in $\mathcal{F}$. If $B$ is paracompact, then the Fiber Bundle Construction lemma implies a one-to-one correspondence between $\check{H}^{1}(B;\mathcal{C}_{\text{Homeo}(F)})$ and $\text{Bun}_{F}(B)$, where $\mathcal{C}_{G}$ denotes the sheaf of continuous $G$-valued functions on $B$ for a topological group $G$. That is, for $U\subseteq B$ open, $\mathcal{C}_{G}(U)=\text{Maps}(U,G)$ with multiplication in $G$.
We state the following additional properties of Čech cohomology without proof. For further details, see [3].
@@
    If $B$ has the homotopy type of a CW complex and $G$ is an abelian group, then there is an isomorphism between Čech and cellular cohomology, $\check{H}^{*}(B;\underline{G})\cong H^{*}(B;G)$, where $\underline{G}$ denotes the sheaf of locally constant $G$-valued functions on $B$.
@@
    If $\mathcal{F}$ is a sheaf of abelian groups over a paracompact space $B$, then there is an isomorphism between sheaf and Čech cohomology, $H^{k}(B;\mathcal{F})\cong\check{H}^{k}(B;\mathcal{F})$, for all $k\geq 0$.
@@
    If $\varphi:G\to H$ is a continuous homomorphism of topological groups, there is an induced cochain map $\varphi^{\#}:\check{C}^{*}(\mathcal{U};\mathcal{C}_{G})\to\check{C}^{*}(\mathcal{U};\mathcal{C}_{H})$ given by $(\varphi^{\#}\Omega)_{\alpha_{0}\cdots\alpha_{k}}=\varphi(\Omega_{\alpha_{0}\cdots\alpha_{k}})$. Moreover, if $\varphi$ is a homotopy equivalence, then $\varphi^{\#}$ induces an isomorphism in cohomology.
  6. 6.
    If $F\to G\to H$ is a central extension of topological groups, there is an induced short exact sequence $\mathcal{C}_{F}\to\mathcal{C}_{G}\to\mathcal{C}_{H}$ of sheaves, which in turn gives rise to an exact sequence in Čech cohomology. Moreover, if $F$ is abelian, there is an induced connecting homomorphism $\Delta:\check{H}^{1}(B;\mathcal{C}_{H})\to\check{H}^{2}(B;\mathcal{C}_{F})$ which yields a long exact sequence.
@@
and $\pi^{\prime}:f^{*}E\to B^{\prime}$ is given by $\pi^{\prime}(b^{\prime},e)=b^{\prime}$. Note that if $\xi$ has fiber $F$, then so does $f^{*}\xi$. Moreover, one can show that homotopic maps yield isomorphic pullback bundles.
@@
For any paracompact space $B$, there is a natural bijective correspondence between $[B,BG]$, the set of homotopy classes of maps from $B$ to $BG$, and $\text{Prin}_{G}(B)$. The correspondence is given by $f]\mapsto[f^{*}\xi_{G}]$, where $\xi_{G}$ is a universal $G$-bundle. Given any $\xi]\in\text{Prin}_{G}(B)$, a map $f:B\to BG$ such that $\xi\cong f^{*}\xi_{G}$ is called a classifying map for $\xi$.
Lastly, we note that the classifying space construction can be made functorial in the following sense: if $\varphi:G\to H$ is a continuous homomorphism of topological groups, there is an induced map $B\varphi:BG\to BH$ with the property that if $f:B\to BG$ classifies a principal $G$-bundle $\xi$, then $B\varphi\circ f:B\to BH$ classifies $\varphi^{*}\xi$.
@@ 2.2 Cohomology With Local Coefficients
We provide a brief overview of cohomology with local coefficients which we use in the classification of circle bundles. For additional details, see [8, Chapter V] or [10, Section 3.I].
@@
Given a topological space $B$ and an abelian group $G$, a local system for $B$ with coefficients in $G$ is a locally constant sheaf $\mathcal{L}$ with stalk $G$.
Note that if $B$ is path-connected and has a universal cover $p:\widetilde{B}\to B$, then a local system on $B$ can equivalently be defined as a sheaf $\mathcal{L}$ with stalk $G$ such that the pullback $p^{*}\mathcal{L}$ is a constant sheaf on $\widetilde{B}$. To see this, note that if $\mathcal{L}$ is a locally-constant sheaf, then $p^{*}\mathcal{L}$ is also locally-constant, but a locally-constant sheaf on a simply-connected space is constant. Conversely, if $\mathcal{L}$ is a sheaf on $B$ and $p^{*}\mathcal{L}$ is constant, then clearly $\mathcal{L}$ is locally-constant.
We will primarily be interested in the case where $G$ is a discrete group and $B$ a path-connected paracompact space. In this setting, a local system on $B$ admits several equivalent characterizations:
@@
Suppose $G$ is a discrete group and $B$ is a path-connected, locally path-connected and semilocally simply connected space. Endow the automorphism group $\text{Aut}(G)$ with the discrete topology, and let $\underline{\text{Aut}(G)}$ denote the associated locally-constant sheaf on $B$. There are bijective correspondences between:
@@
    Isomorphism classes of local systems on $B$ with stalk $G$,
@@
    Isomorphism classes of flat fiber bundles on $B$ with fiber $G$ and structure group $\text{Aut}(G)$,
@@
    Conjugacy classes of homomorphisms $\rho:\pi_{1}(B)\to\text{Aut}(G)$ (called monodromy representations),
@@
    Cohomology classes $\omega]$ in the Čech cohomology set $\check{H}^{1}(B;\underline{\text{Aut}(G)})$.
@@
1 $\leftrightarrow$ 2: Suppose $\mathcal{L}$ is a local system over $B$ with stalk $G$. Choose an open cover $\mathcal{U}=\{U_{j}\}_{j\in J}$ for $B$ and isomorphisms $\{\phi_{j}:\mathcal{L}|_{U_{j}}\to\underline{G}\}_{j\in J}$. Each map $\phi_{j}\circ\phi_{k}^{-1}:\underline{G}\to\underline{G}$ is locally-constant (hence constant) on $\underline{G}$, thus determines an element $g_{jk}\in\text{Aut}(G)$. The $g_{jk}$’s satisfy the cocycle condition, so the collection of all these maps can be interpreted as a Čech cocycle $g\in\check{Z}^{1}(\mathcal{U};\underline{\text{Aut}(G)})$. By the Fiber Bundle Construction Lemma, $g$ determines a flat fiber bundle with fiber $G$ and structure group $\text{Aut}(G)$ (and furthermore, the isomorphism class of the bundle does not depend on the choice of trivializing cover). Conversely, given a flat fiber bundle $\xi$ with fiber $G$ and structure group $\text{Aut}(G)$, the sheaf of local sections of $\xi$ is a local system on $B$ with stalk $G$. One can check that these constructions are inverses up to isomorphism.
2 $\leftrightarrow$ 3: Suppose we have a representation $\rho:\pi_{1}(B)\to\text{Aut}(G)$. The universal cover $p:\widetilde{B}\to B$ of $B$ is a principal $\pi_{1}(B)$-bundle, where $\pi_{1}(B)$ acts on $\widetilde{B}$ from the left by deck transformations (and taking inverses yields the corresponding right action). In particular, this bundle is flat since $\pi_{1}(B)$ is a discrete group. Since $\rho$ gives a left $\pi_{1}(B)$-action on $G$ by automorphisms, so we can form the associated bundle $\widetilde{B}\times_{\rho}G\to B$. Observe that if $\rho^{\prime}:\pi_{1}(B)\to\text{Aut}(G)$ is another representation in the same conjugacy class as $\rho$, then by the Fiber Bundle Construction Lemma, the resulting bundles are isomorphic.
Conversely, let $\mathcal{L}$ be a local system on $B$ with stalk $G$. Choose any $b_{0}\in B$ and isomorphism $\mathcal{L}_{b_{0}}\cong G$, as well as a loop $\gamma:I\to B$ based at $b_{0}$. Since $I=[0,1]$ is contractible, the sheaf $\gamma^{*}\mathcal{L}$ is constant, and the restriction maps induce isomorphisms $(\gamma^{*}\mathcal{L})_{0}\to(\gamma^{*}\mathcal{L})(I)\to(\gamma^{*}\mathcal{L})_{1}$. We therefore obtain an induced map $T_{\gamma}\in\text{Aut}(G)$ via our identification of $(\gamma^{*}\mathcal{L})_{1}=(\gamma^{*}\mathcal{L})_{0}=\mathcal{L}_{b_{0}}$ with $G$. One can check that the automorphisms induced by homotopic loops coincide, and that the assignment $\gamma]\longmapsto T_{\gamma}$ defines a representation $\rho:\pi_{1}(B,b_{0})\to\text{Aut}(G)$. Note that if $b_{1}\in B$, there exists a path $\widetilde{\gamma}:[0,1]\to B$ from $b_{0}$ to $b_{1}$, and we have $\pi_{1}(B,b_{1})\cong\pi_{1}(B,b_{0})$ via the assignment $\gamma]\longmapsto[\widetilde{\gamma}*\gamma*\widetilde{\gamma}^{-1}]$. Similarly, different identifications $\mathcal{L}_{b_{0}}\cong G$ yield homomorphisms which differ by conjugation. Thus, an isomorphism class of local systems with stalk $G$ determines a representation $\rho:\pi_{1}(B)\to\text{Aut}(G)$ up to conjugacy.
@@
A local system with stalk $\mathbb{Z}$ is called an orientation sheaf. Given a path-connected, semilocally simply connected space $B$, an open cover $\mathcal{U}$ of $B$ and a cocycle $\omega\in\check{Z}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})$ (where $\mathbb{Z}_{2}=\{1,-1\}\cong\mathrm{Aut}(\mathbb{Z}))$, we will denote the orientation sheaf corresponding to $\omega$ by $\mathbb{Z}_{\omega}$.
Remark 2.7.
If $B$ is a smooth manifold, the orientation sheaf corresponding to any representative of the Stiefel-Whitney class $w_{1}(TB)\in\check{H}^{1}(B;\underline{\mathbb{Z}_{2}})$ of its tangent bundle, is called an orientation sheaf for the manifold. By Proposition 2.5, $B$ has an orientation sheaf $\mathcal{O}$, unique up to isomorphism.
Let $\mathcal{U}$ be an open cover of a path-connected, locally path-connected, semilocally simply connected space $B$, and let $\mathbb{Z}_{\omega}$ denote the orientation sheaf on $B$ associated with a cocycle $\omega\in\check{Z}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})$. Proposition 2.5 implies we can think of $\mathbb{Z}_{\omega}$ as the sheaf of locally-constant sections of an associated bundle $B_{\omega}\times_{\mathbb{Z}_{2}}\mathbb{Z}$, where $B_{\omega}$ is a double cover of $B$ and $B_{\omega}\to B$ is a principal $\mathbb{Z}_{2}$-bundle. More generally, given a topological group $G$ with a $\mathbb{Z}_{2}$-action, we will be interested in sheaves of sections of associated bundles of the form $B_{\omega}\times_{\mathbb{Z}_{2}}G$:
@@
$\mathcal{P}_{G,\omega}(V)=\{\text{families }\{f_{j}\in\mathcal{C}_{G}(U_{j}\cap V)\}:\ f_{j}|_{U_{jk}\cap V}=\omega_{jk}\cdot f_{k}|_{U_{jk}\cap V}\}$
The sheaf $\mathcal{C}_{G,\omega}$ of continuous $G$-valued functions on $B$ twisted by $\omega$ is the sheafification of $\mathcal{P}_{G,\omega}$. Note that $\mathcal{C}_{G,\omega}$ is equivalently defined as the sheaf of sections of the associated bundle $B_{\omega}\times_{\mathbb{Z}_{2}}G$.
Notation: If $G$ is a discrete group, we will write $G_{\omega}$ instead of $\mathcal{C}_{G,\omega}$.
@@
Let $\{f_{j}\in\mathcal{C}_{G}(U_{j})\}_{j\in J}$ such that $f_{j}|_{U_{jk}}=\omega_{jk}\cdot f_{k}|_{U_{jk}}$ whenever $U_{j}\cap U_{k}$ is nonempty, and let $\{f^{\prime}_{j}=\phi_{j}\cdot f_{j}\in\mathcal{C}_{G}(U_{j})\}_{j\in J}$. Since $\omega=\omega^{\prime}\cdot\delta^{0}\phi$, we have
$\tilde{f}_{j}|_{U_{jk}}=\phi_{j}\cdot f_{j}|_{U_{jk}}=(\phi_{j}\omega_{jk})\cdot f_{k}|_{U_{jk}}=\phi_{k}\cdot f_{k}|_{U_{jk}}=\tilde{f}_{k}|_{U_{jk}}$
@@
One can also define a notion of singular (co)homology with local coefficients (in fact, any (co)homology theory can be promoted to a theory with local coefficients). The details are not important for us, but we note that when $B$ is paracompact, the singular and Čech cohomology groups with coefficients in $\mathcal{L}$ are isomorphic. We also have the following generalization of Poincaré duality [31, Theorem 2.4]:
@@
Let $M$ be a closed connected $n$-manifold, and let $\mathcal{L}$ be a local system of abelian groups. Then $H^{k}(M;\mathcal{L})\cong H_{n-k}(M;\mathcal{L}\otimes\mathcal{O})$ for all $k$, where $\mathcal{O}$ denotes the orientation sheaf of $M$.